Use the elimination method to solve the given system of equations. Please show all work
8x + 11y = 20
5x - 11y = -59

Found 2 solutions by jim_thompson5910, Alan3354:
Answer by jim_thompson5910(35256)

Same idea as last time. Add the equations


8x + 11y = 20
5x - 11y = -59
---------------
13x + 0y = -39


So if 13x = -39, then x = -3 (divide both sides by 13)


I'll let you finish up and find y.

Answer by Alan3354(69443)

8x + 11y = 20
5x - 11y = -59
----------------------- Add them